jigglypuff314 (jigglypuff314):

Let y = f(x) be a particular solution to the differential equation dy/dx = 1 + y^2 - 2x - 2xy^2 with f(2) = 0 Write an equation for the line tangent to the graph of y=f(x) at x=2

jigglypuff314 (jigglypuff314):

@sourwing ? :3

jigglypuff314 (jigglypuff314):

random guess-> y = -2(x-2) ? :P

OpenStudy (anonymous):

did you get dy/dx = (1+y^2) (1 - 2x) ?

jigglypuff314 (jigglypuff314):

yes... for the "Find the particular solution y = f(x) to the differential equation with the initial condtion f(2) = 0. (Hint: Factor out by grouping)" part

OpenStudy (anonymous):

ok then dy/(1+y^2) = (1-2x) dx arctan(y) = x - x^2 + C

OpenStudy (anonymous):

oh wait, actually you don't have to solve the differential equation dy/dx is a slope in terms of x and y. Given (x,y) = (2,0) then dy/dx = (1 + 0^2) (1 - 2(2)) = -3 using point slope form y - 0 = -3(x-2)

jigglypuff314 (jigglypuff314):

oh right okay :) thanks

OpenStudy (anonymous):

np

jigglypuff314 (jigglypuff314):

it's a multipart problem tho how would I do the "Find the particular solution y = f(x) to the differential equation with the initial condtion f(2) = 0. (Hint: Factor out by grouping)" part?

OpenStudy (anonymous):

arctan(y) = x - x^2 + C, and (x,y) = (2,0). Plug the point in and solve for C

jigglypuff314 (jigglypuff314):

alright :) thanks again ^_^

jigglypuff314 (jigglypuff314):

C = 2?

OpenStudy (anonymous):

oh yeah it's 2 :D

jigglypuff314 (jigglypuff314):

lol ok xD so I would write it as arctany = -x^2 + x + 2 ? or try to isolate y somehow?

OpenStudy (anonymous):

if you want to

OpenStudy (anonymous):

it looks like the question wants you to solve for y since it said y = f(x)